Find the value of 4y+2 given that -11y-9=2. Simplify your answer as much as possible.
Answer by stanbon(75887)

Find the value of 4y+2 given that -11y-9=2. Simplify your answer as much as possible.
-------------------
-11y - 9 = 2
-11y = 11
y = -1
-----------
4y = -4
4y+2 = -4+2
4y+2 = -2
===============
Cheers,
Stan H.